OpenStudy (vegamontiel):

If I have a 400lbg weight , and I need to pull it 1 meter , and it takes me 1.5 seconds to pull it . How much force am I using ? ... Do I need to use kinetic energy formula ? Or just the regular force =ma/m^2??

OpenStudy (ljetibo):

that's a weird looking "regular" force. The one I know about looks like F=ma. I suggest you take a peak at momentum and/or impulse to solve this task. $$F \Delta t = \Delta p$$

OpenStudy (vegamontiel):

So what equation would you suggest I use ?? Because I have all the numbers and weights . I'm just not sure what equation to use .

OpenStudy (vegamontiel):

Because I think I also need to include natural force of gravity right ??? An direction ?

OpenStudy (ljetibo):

Let's say you've raised the weight 1m up. Then the change in gravitational potential energy is Epot = mgh = 181.436948kg * 9.81 * 1 = 1778.0820904000002 Joules that is to say that you've done ~1780Joules of work. You've done it in 1.5 seconds: P=W/t = F*h/t = 1185.38806027Watts so it's been done by using 1185.4 Watts of Power. On average you've pulled it with the force of: F = t*1185/h = 1.5*1185/1 = 1177N which means that Ft = mv -> v = Ft/m = 9.7272727 m/s or that you've accelerated this object from 0 to ~10m/s in a span of a second and a half.